OpenStudy (anonymous):

A boy throws a football horizontally from a window that is 2.8 m above the ground. If he throws the ball with a speed of 12 m/s, how far from the window does the ball land?

OpenStudy (shamim):

Given h=2.8 m u=0 t=? g=9.8m/s

OpenStudy (shamim):

Required formula is h=ut+(1/2)gt^2 2.8=0*t+(1/2)*9.8*t^2 t=?

OpenStudy (shamim):

Now horizontal distance s=vt=12*t=?

OpenStudy (shamim):

response plz!!!!